# NPL-MVPS: Neural Point-Light Multi-View Photometric Stereo

### v2

@@ 1 Introduction
Photometric Stereo (PS) is a long standing and very important problem in the field of Computer Vision. While early PS works [43, 14, 31, 16] primarily tackled the estimation of normals from a single view image, the value of PS was unlocked by binocular [21, 7, 42, 28] and multi-view [8, 50, 38, 32, 24, 44] stereo setups as it allowed for accurate recovery of shape and not only normals. This, in turn, opened many applications such as general 3D reconstruction, novel-view rendering, relighting and material editing [44], as well as robot interaction, quality control in manufacturing and industrial conveyor belt scanning.
@@
Figure 1: In this figure we demonstrate the fragility of relying mainly on estimated normals (using [12]) for deep learning based sparse multi-view photometric stereo when 6 views out of 20 and 6 light sources out of 96 available are used. The first column shows the ground truth of Buddha object from dense MVPS DiLiGenT-MV benchmark. The following two columns show the estimated normals and corresponding error map using only 6 out of 96 lights available (view 1 mean average normal error is $9.9^{\circ}$, saturated red color corresponds to $5^{\circ}$ error). Using such normals leads to a large error in reconstruction error using our method when pixel intensities are not leveraged (0.51mm) and previous SOTA dense MVPS method Supernormal [4] (0.67mm). If pixel intensities are used along with estimated normals (column 6) a significantly smaller error of 0.35mm is achieved. The final two columns show the error maps of estimated shapes when all available views and lights are used where Supernormal [4] achieves a similar reconstruction error as our method (0.21mm vs 0.19mm). Similar dynamics apply to other DiLiGenT objects as shown in Tables 2 and 3, strongly motivating for explicit pixel intensity modeling in MVPS methods. Note here the errors are computed as Chamfer distance while the visualisation only shows errors from reconstruction to ground truth mesh for each ground truth mesh surface point. Note dark red corresponds to 1mm error in the shape error illustrations (columns 3-8).
@@
Along with increasing the number of views Photometric Stereo undergone another important change by moving from classical non-linear optimisation enabled inverse graphics approaches (for single view [13], binocular [21], multi-view [32, 24]) to neural network (e.g. [14, 30]) and in particular neural shape representation enabled inverse graphics approaches (for single view [11], binocular [28] and multi-view [44, 4]). However, despite the latter methods, especially [4], achieving impressive accuracy on DiLiGenT-MV [24] benchmark their approach to MVPS is somewhat incomplete as they do not attempt to directly explain (and learn to match) observed pixel-wise intensities. In particular,  [4] does not explicitly use image intensities to optimise for shape and it is fully reliant on per-view normal maps. Whereas, PS-NeRF [44] only uses average intensity during the surface optimisation stage and thus leaves most of the photometric information unused. It is important to note that Brahimi et. al. [1] , attempts to re-render the images however does not model cast shadows and uses the simplified Dinsey BRDF [3] which may not model all materials accurately.
@@
In this work we provide the first neural multi-view photometric stereo approach which fully leverages the availability of pixel intensity information for estimating 3D shape from Photometric Stereo images (see Figure 5). We achieve this by explicitly modeling the incident light from point light sources to leverage intensity based shape optimisation over purely normal driven shape optimisation [44, 4] which is fragile to incorrectly estimated normals especially in cases of very few available lights as shown Figure 1 and Section 5.
@@ 2 Related Work
Neural surfaces. Recently, neural surface approaches have became very popular in tacking the 3D reconstruction problem. Early approaches include NeRF [36] and its first extensions to neural SDF parameterisations [47, 46]. The first methods which used neural SDFs specifically for the multi-view PS problem include [20, 18, 19, 49]. However, contrary to the direction of the neural inverse rendering literature, these approaches do not attempt to re-render the original photometric stereo images but rather some 2D derivates of the images such as normals or albedo maps. For example, Supernormal [4] only renders normal maps but achieves very fast training speed though patch parameterisation, as well as the use of the NERFACC [25] framework. PS-NeRF [44] renders normal and average intensity maps whereas RNb-Neus [2], uses normal and albedo maps to render virtual light images. Thus, all these methods are significantly reliant on single view PS networks and have no way to circumvent noisy estimates that are likely to happen in case of sparse lights and number of views as demonstrated in Figure 1.
@@ 3 Method
Figure 2: Schematic of our overall method. Single view PS is used to obtain normal maps. Training the SDF with normal loss obtains a rough surface which is then refined with full volumetric rendering. The rendering procedure is explained in Figure 3. The second row also shows the GT and render images (as grayscale), the rendering error (with red =0.1) as well as the computed shadow map.
This section describes our method for solving the point light multi-view photometer stereo. A high level overview is also shown in Figure 2. Our method is primarily an inverse neural rendering method. Section 3.1 describes the assumed irradiance model. Sections 3.2 and 3.3 describe the underlying neural surface parametrisation and its initialisation, respectively. Section 3.4 describes training losses used.
@@ 3.1 Irradiance equation
Note, here $B(.)$ is assumed to be a general BRDF, $s_{m}\in\{0,1\}$ is an indicator variable to account for cast shadows that completely block direct reflectance. Also note, we assume that indirect reflectance (i.e. self reflections and ambient light) are negligible and can be ignored.
@@ 3.2 Neural SDF
Ray sampling. We follow a volumetric sampling and rendering method similar to VolSDF [46] where the neural SDF is queried in multiple samples on outgoing rays from each image foreground pixel. For each pixel an estimate of the depth is available (initialised from single view PS and occasionally updated during training) and thus most of the samples are concentrated around that depth. However, to allow the surface to evolve and to minimse free space artefacts, additional samples are also sampled in a wider depth range. Following VOLSDF [46], we use the Laplace density function to convert from SDF values $d$ to transparency $t$ as follows:
@@
Finally, alpha blending is used to accumulate depth, normals and rendered intensity at each ray using the standard approach, i.e. transparecny $\alpha=\text{exp}(-t\delta r)$ with $\delta r$ being the distance along the ray and so the surface-ray intersection point $\boldsymbol{\mathbf{X}}_{I}$ is computed as a weighted sum of ray samples $\boldsymbol{\mathbf{X}}_{I}=\sum_{i}(w_{i}\boldsymbol{\mathbf{X_{r}}}_{i})$, with the sample weights $w_{i}$ corresponding to the accumulated opacity. Note that intersection points $\boldsymbol{\mathbf{X}}_{I}$ are further used to compute cast shadows but are not directly rendered. Similar averaging is used to obtain the rendered intensity ($i_{I}=\sum_{i}(w_{i}i_{i})$) and surface normal ($\boldsymbol{\mathbf{N}}_{I}=\sum_{i}(w_{i}\boldsymbol{\mathbf{N_{r}}}_{i})$) both of which are used to compute losses. The ray sampling process is further explained in Figure 3.
Learned BRDF. We follow the approach of [29] where the BRDF is also parameterised as another SIREN network and thus completely learned from the data. This assumes that the material properties are uniform around the target scene except for a scalar albedo variation. We emphasise that we chose to perform grayscale intensity rendering instead of full RGB rendering as this is expected to minimise the synthetic to real gap. Indeed, real RGB images are usually acquired with demosaicing of single intensity values and this procedure is usually optimised to best recover intensity not colour (e.g see [9] used by OpenCV).
To minimise over-fitting, the material BRDF network receives as input only the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{V}}$. In addition, to simplify the learning problem, we follow the principles described in the MERL real material database [34]. To achieve this, the half vector $\boldsymbol{\mathbf{H}}=\frac{\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}}{|\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}|}$ is first computed and the input to the network is the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{H}}$. Finally, we note that the final activation of SIREN part of the BRDF network is exponential and there is a post multiplication with an $\boldsymbol{\mathbf{N}}\cdot\boldsymbol{\mathbf{L}}$ factor so that the BRDF network learns a multiplicative factor over the diffuse reflectance, parameterised as follows:
@@
Albedo. The scalar albedo $\rho$ is learned with another SIREN network which is queried for every sample point.
Shadow estimation. Handling cast shadows is imperative for any successful photometric stereo method. To estimate cast shadows for a ray-surface intersection point $\boldsymbol{\mathbf{X}}_{I}$, we raytrace from that point to the light source following the direction of the lighting vectors $\boldsymbol{\mathbf{L}}_{m}$ computed above. For each ray we take 16 random samples $j$ $\in[2\text{mm},50\text{mm}]$. For all these points, we query the SDF network and accumulate opacity following the same volumetric rendering procedure, i.e $s=\prod_{j}(1-\alpha_{j})$. We note that this shadow computation procedure is very computationally expensive therefore it cannot be computed for all the rendered points. Instead, we only compute it for the intersection point of each ray ($\boldsymbol{\mathbf{X}}_{I}$ is Figure 3) and assume it is the same for all other ray samples.
@@ 3.3 Initialisation
As it is standard in MVPS approaches, e.g. [44], we can use single view PS (at each view) in order to obtain normal and depth estimates. We start by computing per view normal maps using the state-of-the-art PS normal estimation network [12] and also use numerical integration [39] in order to get approximate depth maps. The depth maps do need to be accurate as they are only used to initialise the ray search spaces. The normal maps are used in order to provided an additional training signal.
@@
Final surface calculation. After the optimisation is completed, the SDF network can be sampled in a regular grid of points and a triangle mesh surface can be recovered using the standard marching cubes [33] algorithm. For these revered vetrices, the albedo network can be queried in order to obtained a textured reconstruction.
Figure 4: Qualitative visualisation of re-rendering and rendering errors for synthetic and real data. The scaling of the error map sets red to 0.1. We note that most of the error is concentrated on the middle of concavities as self reflection is not modeled.
@@ 3.4 Losses
Silhouette loss. We also apply binary cross entropy loss to the predicted silhuettes to the ground truth ones. We generate predicted silhuettes with the total accumulated opacity $\alpha_{total}=\prod_{j}(1-\alpha_{j})$ (which should be 1 for foreground and 0 for background points), $L_{sil}=\text{BCE}(\alpha_{total},mask)$. We consider a 10 pixel band outside the provided segmentation masks for background points. The relative weighting of this loss is 10.
@@
Curvature regulariser. To minimise floater artefacts (especially on the inside of objects) and encourage the optimisation to recover the minimum surface, it important to include some curvature regulariser encourage smoothness on the volumetric normals (i.e. SDF gradients). Computing exact analytic curvatures (via auto-differentiation) has a high computational cost an is not really required, as the objective is to only use them as a regulariser. Instead, drawing inspiration from SuperNormal [4], we use the ray samples $\boldsymbol{\mathbf{X}}_{i}$ and compute finite differences along the ray as: $curv(\boldsymbol{\mathbf{X}}_{i})\approx\frac{||\boldsymbol{\mathbf{N}}_{i}-\boldsymbol{\mathbf{N}}_{i+1}||}{||\boldsymbol{\mathbf{X}}_{i}-\boldsymbol{\mathbf{X}}_{i+1}||}$. We note that an exact curvature would require finite differences along all 3-axis, but for regularisation purposes this definition is adequate, and comes with no additional SIREN queries. Note that NeILF++ [48] also includes a similar regulariser. The relative weighting of this loss is 1.
@@ 4 Experiment Setup
I 0.13 0.29 0.07 0.20 0.11 0.16
N + I 0.11 0.17 0.04 0.18 0.11 0.12
Table 1: Ablation study of our method on a synthetic replica of DiLiGenT-MV [24] benchmark. We first compute our method performance using ground truth normals in order to highlight potential issues with real DiLiGenT-MV [24] benchmark (first two rows of Table 2) where recovered shape is significantly less accurate both for our and Supernormal [4] methods (0.05mm vs 0.11mm). We also include comparison of the 3 versions of our method named N, I and N+I namely normal loss only (no rendering), rendering loss only (no normals) and combined.The combined approach achieves the best error and particularly note that it outperforms both other configurations on Bear and Cow objects indicating the the combined approach is better than a simple interpolation between the two.
This section explains the datasets used for evaluation of our method along with the evaluation metric and competitors.

### v3

@@ 1 Introduction
Photometric Stereo (PS) is a long standing and important problem in the field of Computer Vision. While early PS works [46, 33, 15, 31, 17] primarily tackled the estimation of normals from single view images, the value of PS was unlocked by binocular [22, 8, 44, 29] and multi-view [9, 53, 40, 32, 25, 47] stereo setups as it allowed for accurate recovery of shape and not only normals. This, in turn, opened many applications such as general 3D reconstruction, novel-view rendering, relighting and material editing [47], as well as robot interaction, quality control in manufacturing and industrial conveyor belt scanning.
@@
Figure 1: In this figure we demonstrate the fragility of relying mainly on estimated normals (using [13]) for deep learning based sparse multi-view photometric stereo when 6 views out of 20 and 6 light sources out of 96 available are used. The first column shows the ground truth of Buddha object from dense MVPS DiLiGenT-MV benchmark. The following two columns show the estimated normals and corresponding error maps using only 6 out of 96 lights available (view 1 mean average normal error is $9.9^{\circ}$, saturated red color corresponds to $5^{\circ}$ error). Using such normals leads to a large reconstruction error using our method when pixel intensities are not leveraged (0.51mm) and previous SOTA dense MVPS method Supernormal [4] (0.67mm). If pixel intensities are used along with estimated normals (column 6) a significantly smaller error of 0.35mm is achieved. The final two columns show the error maps of estimated shapes when all available views and lights are used. In this setting Supernormal [4] achieves a similar reconstruction error as our method (0.21mm vs 0.19mm). Similar dynamics apply to other DiLiGenT-MV objects as shown in Tables 2 and 3, strongly motivating for explicit pixel intensity modeling in MVPS methods. Note here the errors are computed as Chamfer distance while the visualisation only shows errors from reconstruction to ground truth mesh for each reconstructed mesh surface point. Note dark red corresponds to $\geq 1mm$ error in the shape error illustrations (columns 3-8).
@@
Along with increasing the number of views Photometric Stereo undergone another important change by moving from classical non-linear optimisation enabled inverse graphics approaches (for single view [14], binocular [22], multi-view [32, 25]) to neural network (e.g. [15, 30]) and in particular neural shape representation enabled inverse graphics approaches (for single view [12], binocular [29] and multi-view [47, 4]). However, despite the latter methods, especially [4], achieving impressive accuracy on DiLiGenT-MV [25] benchmark their approach to MVPS is somewhat incomplete as they do not attempt to directly explain (and learn to match) observed pixel-wise intensities. In particular,  [4] does not explicitly use image intensities to optimise for shape and is fully reliant on per-view normal maps. Whereas, PS-NeRF [47] only uses average intensity during the surface optimisation stage and thus leaves most of the photometric information unused. It is important to note that Brahimi et. al. [1] , attempts to re-render the images however does not model cast shadows and uses the simplified Dinsey BRDF [3] which may not model all materials accurately.
@@
In this work we provide the first neural multi-view photometric stereo approach which fully leverages the availability of pixel intensity information for estimating 3D shape from Photometric Stereo images (see Figure 5). We achieve this by explicitly modeling the incident light from point light sources to leverage intensity based shape optimisation over purely normal driven shape optimisation [47, 4] which is fragile to incorrectly estimated normals especially in cases of very few available lights as shown in Figure 1 and Section 5.
@@ 2 Related Work
Neural surfaces. Recently, neural surface approaches have became very popular in tacking the 3D reconstruction problem. Early approaches include NeRF [38] and its first extensions to neural SDF parameterisations [50, 49, 45]. The first methods which used neural SDFs specifically for the multi-view PS problem include [21, 19, 20, 52]. However, contrary to the direction of the neural inverse rendering literature, these approaches do not attempt to re-render the original photometric stereo images but rather some 2D derivates of the images such as normals or albedo maps. For example, Supernormal [4] only renders normal maps but achieves very fast training speed though patch parameterisation, as well as the use of the NERFACC [26] framework. PS-NeRF [47] renders normal and average intensity maps whereas RNb-Neus [2], uses normal and albedo maps to render virtual light images. Thus, all these methods are reliant on single view PS networks and have no way to circumvent noisy estimates that are likely to happen in case of sparse lights and number of views as demonstrated in Figure 1.
@@ 3 Method
Figure 2: Schematic of our overall method. Single view PS is used to obtain normal maps. Training the SDF with normal and silhouette loss (for 3 epochs only, see Section 3.3) obtains a rough surface which is then refined with full volumetric rendering, explained in Figure 3. The second row also shows the GT and render images (as grayscale), the rendering error (with red $\geq 0.1$) as well as the computed shadow map.
This section describes our method for solving the point light multi-view photometric stereo. A high level overview is also shown in Figure 2. Our method is primarily an inverse neural rendering method. Section 3.1 describes the assumed irradiance model. Sections 3.2 and 3.3 describe the underlying neural surface parametrisation and its initialisation, respectively. Section 3.4 describes training losses used.
@@ 3.1 Irradiance equation
Note, here $B(.)$ is assumed to be a general BRDF, $s_{m}\in\{0,1\}$ is an indicator variable to account for cast shadows that completely block direct reflectance. We assume that indirect reflectance (i.e. self reflections and ambient light) are negligible and can be ignored. Also note that analytic BRDF models like [7, 3] often (but not always [36]) separate diffuse and specular components and may include separate albedos. However, a completely learned BRDF as proposed in Section 3.2 does not need to follow this structure.
@@ 3.2 Neural SDF
Ray sampling. We follow a volumetric sampling and rendering method similar to VolSDF [49] where the neural SDF is queried in multiple samples on outgoing rays from each image foreground pixel. For each pixel an estimate of the depth is available (initialised from single view PS and occasionally updated during training) and thus most of the samples are concentrated around that depth. However, to allow the surface to evolve and to minimise free space artefacts, additional samples are also sampled in a wider depth range. Following VolSDF [49], we use the Laplace density function to convert from SDF values $d$ to density $t$ as follows:
@@
Finally, alpha blending is used to accumulate depth, normals and rendered intensity at each ray using the standard approach, i.e. transparency $\alpha=\text{exp}(-t\delta r)$ with $\delta r$ being the distance along the ray and so the surface-ray intersection point $\boldsymbol{\mathbf{X}}_{I}$ is computed as a weighted sum of ray samples $\boldsymbol{\mathbf{X}}_{I}=\sum_{i}(w_{i}\boldsymbol{\mathbf{X_{r}}}_{i})$, with the sample weights $w_{i}$ corresponding to the accumulated opacity. Note that intersection points $\boldsymbol{\mathbf{X}}_{I}$ are further used to compute cast shadows but are not directly rendered. Similar averaging is used to obtain the rendered intensities ($i_{I}=\sum_{i}(w_{i}i_{i})$) and surface normals ($\boldsymbol{\mathbf{N}}_{I}=\sum_{i}(w_{i}\boldsymbol{\mathbf{N_{r}}}_{i})$) both of which are used to compute losses. The ray sampling process is further explained in Figure 3.
Learned BRDF. We follow the approach of [29] where the BRDF is also parameterised as another SIREN network and thus is completely learned from the data. This assumes that the material properties are uniform around the target scene except for a scalar albedo variation. We emphasise that we chose to perform grayscale intensity rendering instead of full RGB rendering as this is expected to minimise the synthetic to real gap. Real RGB images are usually acquired with demosaicing of single intensity values and this procedure is usually optimised to best recover intensity not colour (e.g see [10] used by OpenCV).
To minimise over-fitting, the material BRDF network receives as input only the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{V}}$. In addition, to simplify the learning problem, we follow the principles described in the MERL database [35]. To achieve this, the half vector $\boldsymbol{\mathbf{H}}=\frac{\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}}{|\boldsymbol{\mathbf{L}}+\boldsymbol{\mathbf{V}}|}$ is first computed and the input to the network is the relative angles between $\boldsymbol{\mathbf{N}}$, $\boldsymbol{\mathbf{L}}$ and $\boldsymbol{\mathbf{H}}$. Finally, we note that the final activation of SIREN part of the BRDF network is exponential and there is a post multiplication with an $\boldsymbol{\mathbf{N}}\cdot\boldsymbol{\mathbf{L}}$ factor so that the BRDF network learns a multiplicative factor over the diffuse reflectance, parameterised as follows:
@@
Albedo. The scalar albedo $\rho$ is learned with another SIREN network which is queried for every sample point. We note that having the BRDF network constant throughout the volume and only varying a scalar albedo may sacrifice quality in objects with significantly varied materials, but this does not seem to be the case in the DiLiGenT-MV [25] as shown in Figure 4. We note some competitors like Neuralangelo [27] learn a fully-varied rendering network parameterised by position, normal, lighting and viewing vectors, but this approach is a lot more prone to over-fitting and would struggle to extrapolate the rendering into completely unseen viewing angles, which is not the case for our approach (see Figure 4).
Shadow estimation. To estimate cast shadows for a ray-surface intersection point $\boldsymbol{\mathbf{X}}_{I}$, we raytrace from that point to the light source following the direction of the lighting vectors $\boldsymbol{\mathbf{L}}_{m}$ computed above. For each ray we take 16 random samples $j$ $\in[2\text{mm},50\text{mm}]$. For all these points, we query the SDF network and accumulate opacity following the same volumetric rendering procedure, i.e $s=\prod_{j}(1-\alpha_{j})$. We note that this shadow computation procedure is very computationally expensive therefore it cannot be computed for all the rendered points. Instead, we only compute it for the intersection point of each ray ($\boldsymbol{\mathbf{X}}_{I}$ is Figure 3) and assume it is the same for all other ray samples.
@@ 3.3 Initialisation
As it is standard in MVPS approaches, e.g. [47], we can use single view PS (at each view) in order to obtain normal and depth estimates. We start by computing per view normal maps using the state-of-the-art PS normal estimation network [13] and also use numerical integration [41] in order to get approximate depth maps. The normal maps are used in order to provide an additional training signal.
The depth maps do need to be accurate as they are only used to initialise the ray sampling space. More specifically, classical ray marching (e.g. NeRF [38]) uses fixed near/far planes for each pixel which is inefficient while newer approaches use an occupancy grid (e.g NERFACC [26]) to guide the search space. We opt for a simpler solution where for each pixel, near/far planes are centered around the pixels depth estimate; this is initialised with single view depth estimate and updated every 5 epochs.
@@
Final surface calculation. After the optimisation is completed, the SDF network can be sampled in a regular grid of points and a triangle mesh surface can be recovered using the standard Marching Cubes [34] algorithm. For recovered surface points, the albedo network is queried in order to obtained a textured reconstruction.
Figure 4: Qualitative visualisation of re-rendering and rendering errors for synthetic and real data (left side, top and bottom 3 rows). The scaling of the error map sets red to $\geq 0.1$. We note that most of the error is concentrated on the middle of concavities as self reflection is not modeled. On the right side we see renderings in a novel angle of objects recovered from real data as well as recovered albedo maps.
@@ 3.4 Losses
Silhouette loss. Similar to previous approaches [45, 27], we also apply binary cross entropy loss, $L_{sil}=\text{BCE}(\alpha_{total},mask)$, between the predicted silhuettes and the ground truth ones. We generate predicted silhuettes with the total accumulated opacity $\alpha_{total}=\prod_{j}(1-\alpha_{j})$ (which should be 1 for foreground and 0 for background points), We consider a 10 pixel band outside the provided segmentation masks for background points. The relative weighting of this loss is 10.
@@
Curvature regulariser. To minimise floater artefacts (especially on the inside of objects) and encourage the optimisation to recover the minimum surface, it important to include some curvature regulariser encourage smoothness on the volumetric normals (i.e. SDF gradients). Computing exact analytic curvatures (via auto-differentiation) has a high computational cost an is not really required, as the objective is to only use them as a regulariser. Instead, inspired by SuperNormal [4], we use the ray samples $\boldsymbol{\mathbf{X}}_{i}$ and compute finite differences along the ray as: $curv(\boldsymbol{\mathbf{X}}_{i})\approx\frac{||\boldsymbol{\mathbf{N}}_{i}-\boldsymbol{\mathbf{N}}_{i+1}||}{||\boldsymbol{\mathbf{X}}_{i}-\boldsymbol{\mathbf{X}}_{i+1}||}$. We note that an exact curvature would require finite differences along all 3-axis, but for regularisation purposes this definition is adequate, and comes with no additional SIREN queries. Note that NeILF++ [51] and Neuralangelo [27] also include a similar regulariser. The relative weighting of this loss is 1.
@@ 4 Experiment Setup
I 0.13 0.23 0.06 0.08 0.12 0.12
I-S 0.20 0.27 0.04 0.08 0.18 0.15
N + I 0.11 0.15 0.03 0.07 0.11 0.09
Table 1: Ablation study of our method on a synthetic replica of DiLiGenT-MV [25] benchmark. We first compute our method performance using ground truth normals in order to highlight potential issues with real DiLiGenT-MV [25] benchmark (first two rows of Table 2) where recovered shape is significantly less accurate both for our and Supernormal [4] methods (0.05mm vs 0.11mm). We also include comparison of four versions of our method named (N), (I), (I-S) and (N+I). (N) corresponds to only applying normals loss, where (I) and (I-S) corrspond to only using rendering loss with and without shadows respectively. (N+I) combines all losses. The combined approach achieves the best error and particularly note that it outperforms both other configurations most objects indicating the the combined approach is better than a simple interpolation between the two.
This section describes the datasets as well as the training and evaluation protocol.